OpenStudy (butterflyhope):

Urgent *will medal*: Simplify nPn-1 for any positive integer n. Thanks c:

OpenStudy (mimi_x3):

hmmm is it n*P*n-1?

OpenStudy (butterflyhope):

Yeah :P

OpenStudy (mimi_x3):

so we can rearrange it \(n*P*n-1=n*n*P-1=n^2P-1\) Ok so im assuming thats what u r looking for but not 100% sure